OpenStudy (anonymous):

A stone is thrown vertically upward with a speed of 18.0 . (a)How fast is it moving when it reaches a height of 11.0 ? (b)How long is required to reach this height? For part A) I did: y=11m, g=9.81m/s^2, Vy0= 18 m/s, y0=0, (V^2)y=? V^2 y= V^2 y0- 2g (y-y0) --> V^2 y=18^2m/s-2(9.8)(11m-0m) --> Vy=Square root ( 324m/s-215.6)= 10.4m/s Part B) I started out with: y=y0 + Vy0t - (1/2) gt^2 --> 11= 18t m/s-(1/2) 9.8t^2 --> -11 + 18t- 9.8t^2 Do I apply the quadratic formula? I do not know how to answer this, please help! This is due at midnight :(

OpenStudy (anonymous):

4.9t^2 -18t +11=0 \[\Delta'=81 - 4.9*11 =27.1\] \[\sqrt{\Delta'} =5.2\] \[t1 =\frac{ 9-5.2 }{ 4.9 } = 0.77s\] t2 = 2.9s you can think an other way v = -gt +vo 10.4=-9.8t +18 t = 0.77s

OpenStudy (masumanwar):

when it reaches maximum height the velocity will be zero and then it will falling under gravitational acceleration

OpenStudy (anonymous):

reaches a height=maximum height oh, sorry, i am poor English